# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v2

@@ 6 Path-based interpretation
As shown in earlier work, Van den Hof and Ramaswamy (2020), the spectrum condition for data-informativity, can be guaranteed to hold generically, if particular path-based conditions on the network graph are satisfied. As was shown to hold in a related identifiability studies in Hendrickx et al. (2019), rank conditions on particular transfer functions in a dynamic network can generically be translated into conditions on vertex disjoint path connections between sets of nodes in the network graph. Here we copy the formulation from Van den Hof et al. (2023), replacing $\kappa$ by $\mu^{[j]}$:
Proposition 5 (Van den Hof and Ramaswamy (2020))
The spectrum condition $\Phi_{\mu^{[j]}}\succ 0$ for almost all $\omega$, holds generically if in the graph of the network there exist $dim(\mu^{[j]})$ vertex disjoint paths from the node sets $(x,e)$ to $\mu^{[j]}$, and the signals in $x$ are mutually independent and persistently exciting.
Actually it should be sufficient to require only those $x$ signals to be persistently exciting that serve as starting nodes in the vertex disjoint paths.
Vertex disjoint paths are paths that connect two sets of nodes with no intermediate vertex being passed by more than one path. They can be calculated on the basis of the network graph through a max flow min cut algorithm as the Ford-Fulkerson algorithm.
When specifying the above result for the particular situation treated in Section 4, we obtain the following.
@@
The spectrum condition $\Phi_{\mu^{[j]}}\succ 0$ for almost all $\omega$, holds generically if in the graph of the network there exist $dim(u_{\!{\scriptscriptstyle\mathcal{D}}^{c}})+1$ vertex disjoint paths between the node sets $\{x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}$ and $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$.
Proof Since $u$ and $e$ are independent, the spectrum condition $\Phi_{\mu^{[j]}}(\omega)\succ 0$ can be recast into the expression $\Phi_{u_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}(\omega)\succ 0$. Together with the fact that by assumption $\Phi_{x}(\omega)\succ 0$ for all $\omega$, this directly leads to the result. $\Box$
7 Application to single module identification in dynamic networks
@@ 7.1 Single module data-informativity results
In this section we are going to make the step from the isolated open-loop situation as considered before, to the situation where the considered MIMO/MISO system can be part of a dynamic network, including possible feedback loops. Dynamic networks are interconnections of dynamic systems and can be represented in different forms, one of them being the so-called module representation (Van den Hof et al., 2013), defined by
@@
with $w$ an $L$-dimensional vector, $G^{0}(q)$ a hollow rational transfer function matrix, i.e. with zeros on the diagonal entries, in the time shift operator $q$, i.e. $q^{-1}w(t)=w(t-1)$, $e$ an $L$-dimensional vector of white noise processes, $H^{0}(q)$ the rational disturbance model, and $u(t)=R^{0}\cdot r(t)$, accounting for the effect of measured external excitation signals $r$ on the network, with $R^{0}$ a binary matrix and $r$ a $K$-dimensional vector of excitation signals. A non-zero element $G_{k\ell}^{0}(q)$ in $G^{0}(q)$ is referred to as a module.
It is further assumed that the network is stable and well-posed (Van den Hof et al., 2013). In this paper it is also assumed, for simplicity, that all elements in $G^{0}(q)$ are strictly proper.
When targeting on the identification of a single module $G_{ji}^{0}$ in this network, several approaches are available for arriving at a consistent module estimate. An indirect identification method is suitable when there is a sufficient number of $u$ signals present in the network, see e.g Van den Hof et al. (2013); Dankers et al. (2016); Gevers et al. (2018). When noise signals $e$ are required to provide sufficient excitation in the network for consistent estimation of the target module, a direct method is the prime approach, see e.g. Van den Hof et al. (2013); Dankers et al. (2016); Ramaswamy and Van den Hof (2021). In this paper we will focus on the direct method.
@@
This actually concerns the description of part of the network (25), while the network nodes that are not present in $\mathcal{Y}\cup\mathcal{D}$ have been removed from the description. The noise contribution $\bar{H}^{0}(q)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}(t)$, represents the (correlated) noise that affects the output $w_{\!{\scriptscriptstyle\mathcal{Y}}}$, with $\bar{H}^{0}(q)$ being monic, stable and stably invertible, and $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}$ a vector white noise process. The noise model in (26) is typically the result of a spectral factorization, after manipulating the network equation (25). The matrix $\bar{T}^{0}(q)$ represents the effect of external excitation signals on $w_{\!{\scriptscriptstyle\mathcal{Y}}}$. Note that not all excitation signals that are present in the network will appear in the term $\bar{T}^{0}(q)u(t)$. Excitation signals that enter the network on nodes $w_{\!{\scriptscriptstyle\mathcal{D}}}$ will contribute to $w_{\!{\scriptscriptstyle\mathcal{Y}}}$ through their respective input term $w_{\!{\scriptscriptstyle\mathcal{D}}}$. The matrix $\bar{T}^{0}(q)$ can have both dynamic terms and fixed terms, the latter situation typically occuring when an excitation signal directly enters the network on a node in $w_{\!{\scriptscriptstyle\mathcal{Y}}}$.
The single module identification problem is then formulated as the problem of identifying a single module $\bar{G}_{ji}^{0}(q)$ in (26) on the basis of data $\{w_{\!{\scriptscriptstyle\mathcal{D}}\cup\!{\scriptscriptstyle\mathcal{Y}}}(t),u(t)\}_{t=0,\cdots N-1}$, under the condition that $\bar{G}_{ji}^{0}(q)=G_{ji}^{0}(q)$, i.e. the original module in the network. Estimation is done on the basis of the one-step ahead prediction error
@@
In order to arrive at a consistent estimate (Ljung (1999)) of the target module, a set of conditions has been formulated in previous work for arriving at a consistent module estimate:
@@
    In order to guarantee module invariance, i.e. $\bar{G}_{ji}^{0}(q)=G_{ji}^{0}(q)$, a parallel path and loop (PPL) condition [^1] The PPL condition Dankers et al. (2016) requires that each path in the network from $w_{i}$ to $w_{j}$ excluding the path through $G_{ji}$, and every loop around $w_{j}$, pass through a measured node that is included in the predictor input. is assumed to be satisfied, which influences the selection of predictor inputs $w_{\!{\scriptscriptstyle\mathcal{D}}}$, see Dankers et al. (2016).
@@
    In order to cope with confounding variables [^2] A confounding variable is an unmeasured variable that has paths to both the input and output of an estimation problem Pearl (2000). multivariable noise models can be used to model the correlated disturbances, leading to a selection of an appropriate set of node signals in $w_{\!{\scriptscriptstyle\mathcal{Y}}}$ and $w_{\!{\scriptscriptstyle\mathcal{D}}}$, see Ramaswamy and Van den Hof (2021).
@@
    Conditions on data-informativity, see Van den Hof et al. (2023).
@@
    Technical conditions on the presence of delays in the network, see Ramaswamy and Van den Hof (2021).
@@
Let $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ be defined as a minimum disconnecting set from $w_{i}$ to the other components in $w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$, under the constraint that $i\notin\mathcal{D}^{c}$.
@@
  • $w_{\!{\scriptscriptstyle\mathcal{T}}}$ as the set of nodes defined by $w_{\!{\scriptscriptstyle\mathcal{T}}}:=w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}\backslash\{w_{i}\cup w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\}$, i.e. the set of input nodes to parametrized modules, excluding $w_{i}$, that are not in the disconnecting set.
  • $w_{\!{\scriptscriptstyle\mathcal{F}}}^{[j]}$ as the set of nodes that are inputs to a non-parametrized (known) module in the $j$-th row of the predictor model.
@@
Consider the network equation (25) for $w$ and consider the node sets $w_{i},w_{\!{\scriptscriptstyle\mathcal{T}}},w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ as defined above. Then there exist rational transfer function matrices $T_{s}(q)$ and $R_{s}(q)$ such that $w_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as:
@@
Consider the network (25), predictor model (26)-(27), the signal vector $\kappa^{[j]}$ (31) and the node sets $w_{i},w_{\!{\scriptscriptstyle\mathcal{D}}^{c}},w_{\!{\scriptscriptstyle\mathcal{T}}},x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ as defined above. If
@@
One of the main differences with the situation of Section 4 is that the two components in the vector $\eta^{[j]}$ are no longer independent. Because of possible feedback loops present in the network, the concerned signals can be correlated. An implication of the spectrum condition is then that $w_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})}$ should be persistently excited, through the presence of external signals that exclude $\chi$ and $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, while $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$ should be persistently exciting in itself. This latter condition is due to the fact that these signals are inputs to parametrized terms in the noise model, and therefore cannot be used for exciting the elements in $G$.
The external signals $\chi$ that are excluded from serving as excitation sources for estimating $G_{ji}^{0}(q)$, now also include all $u$-signals that are input to a parametrized module towards the target output $w_{j}$, i.e. the inputs that correspond to parametrized terms in $[\bar{T}_{(}q,\theta)]_{j*}$, (27). Note however that, different from the term $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, the signals in $u^{[j]}$ do not require excitation, as they are not part of the vector signal $\eta^{[j]}$.
@@
Consider the situation of Theorem 8, and consider the sets of external signals, defined by:
  • $e^{\perp\!{\scriptscriptstyle\mathcal{Y}}}$: All $e$ signals in the network that do not have an unknown path to $w_{\!{\scriptscriptstyle\mathcal{Y}}}$, and
@@
Then the data informativity condition (a) in Theorem 8 is generically satisfied if there exist $dim(\mathcal{D}^{c})+1$ vertex disjoint paths from $(e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j})$ to $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, that do not pass through $w_{\!{\scriptscriptstyle\mathcal{T}}}$, and the signals in $(e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j})$ are mutually independent and persistently exciting.
Proof See Appendix $\Box$
The class of external signals $u^{\perp j}$ can alternatively be specified by the set of $u$-signals, that contains $u_{j}$ (the $u$-signal directly entering $w_{j}$) plus the $u$ signals that have a path to $(w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]},w_{\!{\scriptscriptstyle\mathcal{F}}}^{[j]})$ without passing through $w_{j}$.
The result shows that for appropriately exciting $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, we can use all external signals that in the predictor model do not have a parametrized link to the output. External signals that do have such a parametrized link are required for exciting the concerned entries in $\bar{H}(q,\theta)$ and $\bar{T}(q,\theta)$ respectively, and therefore are not available for exciting the node inputs for estimating $G_{ji}^{0}(q)$. From the remaining set of external signals, only those external signals can be used that reach the predictor input signals through the set of nodes $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, without passing through $w_{\!{\scriptscriptstyle\mathcal{T}}}$.
@@
The set $e^{\perp\!{\scriptscriptstyle\mathcal{Y}}}$ as formulated in Proposition 9, can in some specific cases of predictor models, be slightly enlarged. This holds for predictor models that allow for predictor inputs that have a confounding variable with an output signal, a so-called nonempty set $w_{\!{\scriptscriptstyle\mathcal{B}}}$ Ramaswamy and Van den Hof (2021). A further specification of this will be presented elsewhere.

### v3

@@ 7 Path-based interpretation
As shown in earlier work VandenHof&Ramaswamy_CDC:20 , a spectrum condition for data-informativity like (16), can be guaranteed to hold generically, if particular path-based conditions on the network graph are satisfied. As was shown to hold in a related identifiability studies in Hendrickx&Gevers&Bazanella_TAC:19 , rank conditions on particular transfer functions in a dynamic network can generically be translated into conditions on vertex disjoint path connections between sets of nodes in the network graph. Two directed paths are called vertex disjoint if they do not share any vertex, including the starting and ending vertices. They can be calculated on the basis of the network graph through a max-flow min-cut algorithm as the Ford-Fulkerson algorithm. For handling this result we copy the formulation from VandenHof&etal_IFAC:23 , replacing $\kappa$ by $\mu^{[j]}$ from (17):
Proposition 5 (VandenHof&etal_IFAC:23 )
The spectrum condition $\Phi_{\mu^{[j]}}\succ 0$ for almost all $\omega$, holds generically [^3] In line with its use in Hendrickx&Gevers&Bazanella_TAC:19 , this concerns a property that holds for all possible dynamic modules in the network, except possibly for a set of measure 0.if in the graph of the network there exist $dim(\mu^{[j]})$ vertex disjoint paths from the node sets ${\color[rgb]{0,0,0}\{}x,e{\color[rgb]{0,0,0}\}}$ to ${\color[rgb]{0,0,0}\{}\mu^{[j]}{\color[rgb]{0,0,0}\}}$, and the signals in $x$ are mutually independent and persistently exciting.
Actually, it should be sufficient to require only those $x$ signals to be persistently exciting that serve as starting nodes in the vertex disjoint paths. When specifying the above result for the particular situation treated in Section 5, we obtain the following.
@@
The graph-based condition in Proposition 5 is satisfied if there exist $dim(u_{\!{\scriptscriptstyle\mathcal{D}}^{c}})+1$ vertex disjoint paths between the node sets $x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}$ and ${\color[rgb]{0,0,0}\{}u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$.
Proof. Since $u$ and $e$ are independent, the spectrum condition $\Phi_{\mu^{[j]}}(\omega)\succ 0$ can be recast into the expression $\Phi_{u_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}(\omega)\succ 0$. Together with the fact that by assumption $\Phi_{x}(\omega)\succ 0$ for all $\omega$, this directly leads to the result. $\Box$
We can illustrate this result for the example in Section 6. In the situation where all $G$-modules are parametrized it was found that ${\color[rgb]{0,0,0}\{}u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}=\{u_{2}\}$, ${\color[rgb]{0,0,0}\{}u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}=\{u_{1},u_{2}\}$ and $x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}=\{x_{1},x_{2}\}$. Therefore, two vertex disjoint paths between $\{x_{1},x_{2}\}$ and $\{u_{1},u_{2}\}$, are sufficient to guarantee (generic) data-informativity. This condition is satisfied if both $x_{1}$ and $x_{2}$ are present.
8 Single module identification in dynamic networks
@@ 8.1 Single module data-informativity results
In this section we are going to make the step from the isolated open-loop situation as considered before, to the situation where the considered MIMO/MISO system can be part of a dynamic network, including possible feedback loops. Dynamic networks are interconnections of dynamic systems and can be represented in different forms, one of them being the so-called module representation (VandenHof&etal_Autom:13, ), defined by
@@
with $w$ an $L$-dimensional vector, $G^{0}(q)$ a hollow rational transfer function matrix, i.e. with zeros on the diagonal entries, $e$ an $L$-dimensional vector of white noise processes, $H^{0}(q)$ the rational disturbance model, and $u(t)=R^{0}\cdot r(t)$, accounting for the effect of measured external excitation signals $r$ on the network, with $R^{0}$ a binary $L\times K$ matrix and $r$ a $K$-dimensional vector of excitation signals. A non-zero element $G_{k\ell}^{0}(q)$ in $G^{0}(q)$ is referred to as a module.
It is further assumed that the network is stable and well-posed (VandenHof&etal_Autom:13, ). In this paper it is also assumed, for simplicity, that all elements in $G^{0}(q)$ are strictly proper.
When targeting on the identification of a single module $G_{ji}^{0}$ in this network, several approaches are available for arriving at a consistent module estimate. An indirect identification method is suitable when there is a sufficient number of $u$ signals present in the network, see e.g VandenHof&etal_Autom:13 ; Dankers&etal_TAC:16 ; Gevers&etal_sysid:18 . When noise signals $e$ are required to provide sufficient excitation in the network for consistent estimation of the target module, a direct method is the prime approach, see e.g. VandenHof&etal_Autom:13 ; Dankers&etal_TAC:16 ; Ramaswamy&VandenHof_TAC:21 . In this paper we will focus on the direct method.
@@
This actually concerns the description of part of the network (20), where the equations for the nodes outside of $w_{\!{\scriptscriptstyle\mathcal{Y}}}$ have been decoupled and removed, while the network nodes that are not present in $w_{\mathcal{Y}\cup\mathcal{D}}$ have been removed (immersed) from the description. The noise contribution $\bar{H}^{0}(q)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}(t)$, represents the (correlated) noise that affects the output $w_{\!{\scriptscriptstyle\mathcal{Y}}}$, with $\bar{H}^{0}(q)$ being monic, stable and stably invertible, and $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}$ a vector white noise process. In general, the noise model in (21) is the result of a spectral factorization, after manipulating the network equation (20); that is why the noise term $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}$ is typically different from $e_{\!{\scriptscriptstyle\mathcal{Y}}}$. The matrix $\bar{T}^{0}(q)$ represents the effect of external excitation signals on $w_{\!{\scriptscriptstyle\mathcal{Y}}}$. Note that not all excitation signals that are present in the network will appear in the term $\bar{T}^{0}(q)u(t)$. Excitation signals that enter the network on nodes $w_{\!{\scriptscriptstyle\mathcal{D}}}$ will generally contribute to $w_{\!{\scriptscriptstyle\mathcal{Y}}}$ through their respective input term $w_{\!{\scriptscriptstyle\mathcal{D}}}$. The matrix $\bar{T}^{0}(q)$ can have both dynamic terms and fixed terms, the latter situation, e.g., occuring when an excitation signal directly enters the network on a node in $w_{\!{\scriptscriptstyle\mathcal{Y}}}$.
The single module identification problem is then formulated as the problem of identifying a single module $\bar{G}_{ji}^{0}(q)$, defined as the mapping from $w_{i}$ to $w_{j}$ in (21), on the basis of data $\{w_{\!{\scriptscriptstyle\mathcal{D}}\cup\!{\scriptscriptstyle\mathcal{Y}}}(t),u(t)\}_{t=0,\cdots N-1}$, under the condition that $\bar{G}_{ji}^{0}(q)=G_{ji}^{0}(q)$, i.e. the original module in the network. Estimation is done on the basis of the one-step ahead prediction error
@@
In order to arrive at a consistent estimate (Ljung:99 ) of the target module, a set of conditions has been formulated in previous works:
@@
    In order to guarantee module invariance, i.e. $\bar{G}_{ji}^{0}(q)=G_{ji}^{0}(q)$, a parallel path and loop (PPL) condition [^4] The PPL condition Dankers&etal_TAC:16 requires that each path in the network from $w_{i}$ to $w_{j}$ excluding the path through $G_{ji}$, and every loop around $w_{j}$, pass through a measured node that is included in the predictor input. is assumed to be satisfied, which influences the selection of predictor inputs $w_{\!{\scriptscriptstyle\mathcal{D}}}$, see Dankers&etal_TAC:16 .
@@
    In order to cope with confounding variables [^5] A confounding variable is an unmeasured variable that has paths to both the input and output of an estimation problem Pearl:2000 . multivariable noise models can be used to model the correlated disturbances, leading to a selection of an appropriate set of node signals in $w_{\!{\scriptscriptstyle\mathcal{Y}}}$ and $w_{\!{\scriptscriptstyle\mathcal{D}}}$, see Ramaswamy&VandenHof_TAC:21 .
@@
    Conditions on data-informativity, see VandenHof&etal_IFAC:23 .
@@
    Technical conditions on the presence of delays in the network, see Ramaswamy&VandenHof_TAC:21 .
@@
Let ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$ be defined as a minimum disconnecting set from $w_{i}$ to the other components in ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}{\color[rgb]{0,0,0}\}}$, under the constraint that $i\notin\mathcal{D}^{c}$.
@@
[…]
  • $w_{\!{\scriptscriptstyle\mathcal{F}}}^{[j]}$ as the node vector with nodes that are inputs to a non-parametrized (known) module in the $j$-th row of the predictor model.
@@
Consider the network equation (20) for $w$ and consider the node vectors $w_{i},w_{\!{\scriptscriptstyle\mathcal{T}}},w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ as defined above. Then there exist rational transfer function matrices $T_{s}(q)$ and $R_{s}(q)$ such that $w_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as:
@@
Consider the network (20), predictor model (21)-(22), the signal vector $\kappa^{[j]}$ (26) and the node vectors $w_{i},w_{\!{\scriptscriptstyle\mathcal{D}}^{c}},w_{\!{\scriptscriptstyle\mathcal{T}}},x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ as defined above. If
@@
One of the main differences with the situation of Section 5 is that the two components in the vector $\eta^{[j]}$ are no longer independent. Because of possible feedback loops present in the network, the concerned signals can be correlated. An implication of the spectrum condition is that $w_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})}$ should be persistently excited, through the presence of external signals that exclude $\chi$ and $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, while $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$ should be persistently exciting in itself. This latter condition is due to the fact that these signals are inputs to parametrized terms in the noise model, and therefore cannot be used for exciting the elements in $G$.
The external signals $\chi$ that are excluded from serving as excitation sources for estimating $G_{ji}^{0}(q)$, now also include all $u$-signals that are input to a parametrized module towards the target output $w_{j}$, i.e. the inputs that correspond to parametrized terms in $\bar{T}_{(}q,\theta)]_{j*}$, the $j$-th row of $\bar{T}(q,\theta)$ in (22). Note however that, different from the term $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, the signals in $u^{[j]}$ do not require excitation, as they are not part of the vector signal $\eta^{[j]}$.
@@
Consider the situation of Theorem 8, and consider the selections of external signals, defined by:
  • $e^{\perp\!{\scriptscriptstyle\mathcal{Y}}}$: All $e$ signals in the network that do not have an unknown path [^6] An “unknown path” is a path that contains a link with unknown dynamics that is parametrized in the predictor model. to $w_{\!{\scriptscriptstyle\mathcal{Y}}}$, and
@@
Then the data informativity condition (a) in Theorem 8 is generically satisfied if there exist $dim(\mathcal{D}^{c})+1$ vertex disjoint paths from ${\color[rgb]{0,0,0}\{}e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j}{\color[rgb]{0,0,0}\}}$ to ${\color[rgb]{0,0,0}\{}w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$, that do not pass through ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{T}}}{\color[rgb]{0,0,0}\}}$, and the signals in ${\color[rgb]{0,0,0}\{}e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j}{\color[rgb]{0,0,0}\}}$ are mutually independent and persistently exciting.
Proof See Appendix. $\Box$
The result shows that for appropriately exciting $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, we can use all external signals that in the predictor model do not have a parametrized link to the output. External signals that do have such a parametrized link are required for exciting the concerned entries in $\bar{H}(q,\theta)$ and $\bar{T}(q,\theta)$ respectively, and therefore are not available for exciting the node inputs for estimating $G_{ji}^{0}(q)$. From the remaining set of external signals, only those external signals can be used that reach the predictor input signals through the set of nodes ${\color[rgb]{0,0,0}\{}w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$, without passing through ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{T}}}{\color[rgb]{0,0,0}\}}$.
@@
The selection $e^{\perp\!{\scriptscriptstyle\mathcal{Y}}}$ as formulated in Proposition 9, can in some specific cases of predictor models, be slightly enlarged, in particular when the number of $e$ nodes with an unknown path to $e_{\!{\scriptscriptstyle\mathcal{Y}}}$ exceeds $dim(w_{\!{\scriptscriptstyle\mathcal{Y}}})$. This holds for predictor models that allow for predictor inputs that have a confounding variable with an output signal, a so-called nonempty set $w_{\!{\scriptscriptstyle\mathcal{B}}}$ Ramaswamy&VandenHof_TAC:21 . A further specification of this will be presented elsewhere.